OpenStudy (anonymous):

The manufacturer of a fertilizer guarantees that, with the aid of the fertilizer, 65% of planted seeds will germinate. Suppose the manufacturer is correct. If 6 seeds planted with the fertilizer are randomly selected, what is the probability that more than 4 of them germinate? Carry your intermediate computations to at least four decimal places, and round your answer to at least two decimal places.

OpenStudy (anonymous):

that means five or six right? use the binomial for this one

OpenStudy (anonymous):

I want to do it this way: \[\frac{ 4 }{ 6} * \frac{ 65 }{ 100 }\] but im sure that's too easy/wrong ??

OpenStudy (anonymous):

\[P(x=5)=\binom{6}{5}.65^\times .35\\ P(x=6)=.65^6\]

OpenStudy (anonymous):

I'm confused on the "x=5" thing, since 5 isn't part of the equation. (Sorry! Math is my worst subject.)

OpenStudy (anonymous):

forget the \(x=5\) then

OpenStudy (anonymous):

that just means the probability you get five successes

OpenStudy (anonymous):

Just saw your "that means 5 or 6" comment. I get how you did the x=5 portion now.. I am still confused on the formulas for these.

OpenStudy (anonymous):

lets compute the easy one first

OpenStudy (anonymous):

what is the probability you get all six? the probability of getting any one success is \(.65\) and assuming the trials are independent, (which we do) then the probability of getting a success 6 times in a row means got it on the first try AND on the second try AND on the third try AND on the fourth try AND on the fifth try AND on the sixth i.e. \[.65\times.65\times .65\times .65\times .65\]more succinctly written as \(.65^6\)

OpenStudy (anonymous):

i capitalized AND

OpenStudy (anonymous):

because that means multiply

OpenStudy (anonymous):

I'm getting .0754 for that one

OpenStudy (anonymous):

ok fine looks good now what about the probability you get 5 out of 6?

OpenStudy (anonymous):

that one does make sense to me. but for the 5 would we not just do \[.65^{5}\] ?

OpenStudy (anonymous):

which i confused you by writing \(P(x=5)\)

OpenStudy (anonymous):

oh no you need a \(.65^5\) because that is success AND success AND success AND success AND success

OpenStudy (anonymous):

but what about the failure?

OpenStudy (anonymous):

success AND success AND success AND success AND success AND failure

OpenStudy (anonymous):

\[.65^5\times .35\] if they occur in that order but they could occur in any of 6 different ways SSSSSF SSSSFS SSSFSS SSFSSS SFSSSS FSSSSS

OpenStudy (anonymous):

so you have to take \[.65^5\times .35\] and multiply it by \(6\) i.e. \[6\times .65^5\times .35\]

OpenStudy (anonymous):

I'm so sorry to ask this.. where do you get .35 from?

OpenStudy (anonymous):

i will ask you where do you think it came from? hint: success AND success AND success AND success AND success AND \(\color{red}{\text{failure}}\)

OpenStudy (anonymous):

well I tried doing 1/6 but i got .16 (.17 while rounding) Boy i really got myself in over my head with this stats class!

OpenStudy (anonymous):

there is no 1/6 in the problem if the probability of success is .65 then the probability of failure is 1-.65=.35

OpenStudy (anonymous):

your final answer is \[.65^6+6\times .65^5\times .35\]

OpenStudy (anonymous):

I'm getting 0.3191 (if rounding)

OpenStudy (anonymous):

Thank you for explaining the .35 thing btw. It clicked immediately after you said it, i have a habit of over-complicating things in my head.